Sol Summary – April 19th

Crew 297 Sol Summary Report 19-04-2024

Sol: 05

Summary Title: So much universe, and so little time

Author’s name: Pawel Sawicki

Mission Status: Nominal

Sol Activity Summary:

Today we all woke up early to the sounds of Strauss and the smells of freshly made pancakes by Chef Laude. Shortly thereafter the first of two EVAs commenced, where, within an hour of leaving the airlock, Matt S. located the [inert] Nuclear Power System (NPS). After the crew hid the NPS they met all their EVA objectives and returned to the Hab. The second EVA group was not quite as swift as the first in finding the NPS, but nonetheless Pawel was able to locate the NPS within an hour of EVA start. Collecting gamma-ray spectral data was unfortunately cut short during the second EVA due to a more urgent matter: a nonconforming fan on Suit 11, resulting in an expedited return back to the Hab. The day concluded with a spaghetti dinner prepared by Chefs Laude and Storch, report writing, and late-night games.

Look Ahead Plan:

Tomorrow we have an eventful EVA planned with three of the crew’s research topics consolidating into one EVA objective. The crew of four (Sawicki, Storch, Lynch, and Lamm) will first locate the [inert] NPS. Once found, the crew will use Case (which has recently been equipped with a Geiger counter) to approach the [inert] NPS and verify that the unit is safe to approach. During the recovery efforts, continued use of the gamma-ray spectrometer will further characterize the Martian rocks.

Anomalies in work:

Suit 11 (fan power significantly decreased mid-EVA): EVA was abruptly cut short due to low fan rpm, causing pressing concerns for heat exhaustion. Prior to the 1.5 hour EVA, the suit battery read 13.5 V and afterwards read 11.9 V.
GreenHab automated fan: The GreenHab fan powers on when the GH temp reaches 94F, but then shortly thereafter turns off. It power cycles continuously but never reaches its expected rpm.
ScienceDome: We have noticed a foul chemical smell the last few days which has made the ScienceDome unsuitable for any work. This intense odor has assuaged during the current sol. We still continue to monitor closely.
EVA headset #2 not receiving comms when connected with two way radio (radio works fine).
Morning (data from 0900): calm winds (1 mph), chilly (59F), and few cloud coverage
Noon: calm winds (2 mph), warm (69F), and scattered cloud coverage
Afternoon (1800): moderate winds (17 mph), warm (71F), and broken cloud coverage
Crew Physical Status: Sarah is feeling a bit fatigued, but recovering well. The rest of the crew is presently nominal in terms of health.

EVA: The first EVA today was a complete success. The crew was able to locate the NPS and re-hide it for the next crew in great time. While the second EVA crew was able to locate the NPS quite quickly (~25 minutes), there was a fan malfunction within suit #11 that resulted in an aborted EVA and expedited return to the Hab. Sarah had begun exhibiting symptoms of heat exhaustion from the lack of airflow within her helmet.

To loosely paraphrase Artemis by Weir: "[Mars] is mean. [He] doesn’t care why your suit fails. [He] just [cancels your EVA] when it does."`

Reports to be filed:

Sol Summary
Journalist Report
GreenHab Report
Operations Report
EVA report (2 EVAs)
EVA request (1 EVA)
Support Requested: We have requested that the [inert] NPS be placed by Mission Support at a location of their choosing (though reachable by rover on Galileo Road 1104 and within a maximum of 25 minutes of walking) for a search and recovery EVA #8 on 20-04-2024. The NPS will be placed in the rear airlock tonight, 19-04-2024 by 2000. Once the inert NPS is hidden, precise GPS coordinates and a photo are requested to be taken to not lose track of the location. These GPS coordinates are then asked to be sent to Sean. Many thanks for the support provided thus far for the NPS placements!

We have also requested one additional EVA headset, as one is not functioning properly.

Copyright © The Mars Society. All rights reserved. | Main Site